OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (chriss):

\[m=\frac{2-3}{2-6}=\frac{-1}{-4}=\frac{1}{4}\] \[2=\frac{1}{4}2+b\]\[2=\frac{1}{2}+b\] \[\frac{3}{2}=b\] y=mx+b \[y=\frac{1}{4}x+\frac{3}{2}\]